OpenStudy (anonymous):

solve for x. express your answer as an integer or a simplified fraction. -2(4x+4)-5=-2(x+2)+4x

OpenStudy (goldphenoix):

Do the distributive property. Do you know what -2(4x+4) would look like after distributing?

OpenStudy (anonymous):

-8x-8

OpenStudy (goldphenoix):

Good. What about -2(x+2)?

OpenStudy (anonymous):

-2x-4

OpenStudy (goldphenoix):

Good. Now your equation reads: -8x-8-5 = -2x-4+4x Now combine like term. Which terms can we combine?

OpenStudy (anonymous):

-10x-17+4x

OpenStudy (goldphenoix):

Hmm. Lets do it step-by-step. On the left side of the equation: -8x-8-5. Do you know which terms we can combine using only those three terms?

OpenStudy (anonymous):

-13

OpenStudy (goldphenoix):

Great! So yo know that -8 + -5 = -13. So now we have -8x - 13. Do you think we can combine those two terms?

OpenStudy (anonymous):

no

OpenStudy (goldphenoix):

Good! So now our equation reads: -8x - 13 = -2x-4+4x Which terms can we combine for 2x-4+4x ?

OpenStudy (anonymous):

6x-4

OpenStudy (goldphenoix):

Good, so now our equation looks like: -8x - 13 = 6x -4 Do you know what to do next?

OpenStudy (anonymous):

would it be -2x -13=-4

OpenStudy (goldphenoix):

Not quite. What did you do first?

OpenStudy (anonymous):

-14x-13=-4

OpenStudy (goldphenoix):

Alright, that can work. What's the next step?

OpenStudy (anonymous):

add 13 to both sides

OpenStudy (anonymous):

-14x=9

OpenStudy (anonymous):

-9/14

OpenStudy (anonymous):

thats notone of my answers its either -3/10 -3/2 -9/10 or 3/10

OpenStudy (anonymous):

wait i think i did it wrong i just redid it and got -3/10 is that what you got

OpenStudy (goldphenoix):

Hmm. Divide -14 to both side. What do you get?

OpenStudy (anonymous):

.64

OpenStudy (goldphenoix):

-0.64.

OpenStudy (goldphenoix):

So what do you think the answer is?

OpenStudy (goldphenoix):

I have to go.

OpenStudy (anonymous):

the correct answer is -9/10.-5-2 (4+4 x) = 4 x-2 (2+x) -2 4 x-2 4-5 = -2 x-2 2+4 x -8 x-8-5 = -2 x-4+4 x -8x-13=2x-4 -8x-2x=13-4 -10x=9 x=-9/10